# Normality in the square of the Sorgenfrey line

### v1

@@ 2. Q-sets and the Sorgenfrey square
Let $E$ be any subset of $\mathbb{R}$} that admits no uncountable strictly monotone functions (e.g., the set constructed in Theorem 2.8), then $(E[\leq])^{2}$ is normal.
@@
Let $A,B$ be closed in $(E[\leq])^{2}$. Recall the shoelace lemma: to separate $A$ and $B$ it suffices to cover each set by a countable family of open sets whose closures are disjoint from the other closed set. By symmetry, it suffices to find a countable cover $\{U_{n}\}_{n<\omega}$ of $A$ such that $\overline{U_{n}}\cap B=\emptyset$.
Recall we defined $A^{\prime}\subseteq A$ to have the covering property (with respect to $B$) if there is a countable open cover $\{U_{n}\}$ such that $\overline{U_{n}}\cap B=\emptyset$
Clearly, we have:
Lemma 2.10.
If $A^{\prime}\subseteq A$ is Lindelof, then $A^{\prime}$ has the covering property.
For all $z\in B^{c}$ choose $n_{z}\in\mathbb{N}$ such that $U(z,1/n_{z})\cap B=\emptyset$, this is possible since $B$ is Sorgenfrey closed. By letting $A_{1}=A\cap\overline{B}^{E}$ and $A_{2}=A\setminus\overline{B}^{E}$, we have
Observation 1.
$A=A_{1}\cup A_{2}$
Claim 2.11.
$A_{2}$ has the covering property.
Proof.
Note that $A_{2}\subseteq E\setminus cl_{E}(B)=U$ and the latter is Euclidean open. Since $E^{2}$ is hereditarily Lindelöf with respect to the Euclidean topology, $U$ can be written as the countable union of basic open sets $B_{n}$ such that $\overline{B_{n}}\cap B=\emptyset$ for each $n$ so, since the Sorgenfrey topology is finer than the Euclidean, $A_{2}$ has the covering property. ∎
Claim 2.12.
$A_{1}$ has the covering property.
Proof.
We will break down $A_{1}$ into countably many pieces, each of which is the graph of a monotone decreasing function and prove every piece has the covering property. Define
$A_{q}=\{z=(x,y)\in A_{1}:x+y<q<x+y+1/n_{z}\}$
Denote $l_{q}=\{(x,y)\in X^{2}:x+y=q\}$, $A_{q}$ is just the collection of points $z$ which has the straight line $l_{q}$ passing through the interior of its nbhd $U(z,1/n_{z})$. We see that
Observation 2.
$A_{1}=\cup_{q\in\mathbb{Q}}A_{q}$
So now it suffices to show that each $A_{q}$ has the covering property. First note that
Claim 2.13.
(non-increasing property) For all $z_{1}=(x_{1},y_{1}),z_{2}=(x_{2},y_{2})\in A_{q}$, if $x_{1}<x_{2}$, then $y_{2}\leq y_{1}$.
Proof.
For otherwise we have $x_{1}<x_{2}$ and $y_{1}<y_{2}$. Since the line $l_{q}$ passes through both nbhds $U(z_{1},1/n_{z_{1}}),U(z_{2},1/n_{z_{2}})$, this means $z_{2}$ is in the Euclidean interior of $U(z_{1},1/n_{z_{1}})$. However, $\operatorname{int}_{E}U(z_{1},1/n_{z_{1}})$ is a Euclidean open set, and since $z_{2}\in\overline{B}^{E}$, we have that $\operatorname{int}_{E}U(z_{1},1/n_{z_{1}})$ also contains a point from $B$, which is a contradiction, because we chose $n_{z_{1}}$ so that its nbhd $U(z_{1},1/n_{z_{1}})$ is disjoint from $B$. ∎
Claim 2.14.
The sets $M_{1}=\{x\in p_{0}(A_{q}):|p_{0}^{-1}(x)|>1\},M_{2}=\{x\in p_{1}(A_{q}):|p_{1}^{-1}(x)|>1\}$ are countable.
Proof.
For each $x\in M_{1}$, fix $\alpha_{x}<\beta_{x}\in p_{0}^{-1}(x)$. And for any $x_{1}<x_{2}\in M_{1}$, we have $\beta_{x_{2}}\leq\alpha_{x_{1}}$ by the previous claim. And so the intervals $(\alpha_{x_{1}},\beta_{x_{1}})$ and $(\alpha_{x_{2}},\beta_{x_{2}})$ are disjoint. Hence, we have an uncountable set of pairwise disjoint open intervals on $\mathbb{R}$}, contradiction. The proof for $M_{2}$ is the same. ∎
Let $L_{1}=p_{0}^{-1}(M_{1}),L_{2}=p_{1}^{-1}(M_{2})$, and $L=L_{1}\cup L_{2}$, $K=A_{q}\setminus L$, we see that
Observation 3.
$A_{q}=K\cup L$
Claim 2.15.
$L$ has the covering property.
Proof.
We know $L_{1},L_{2}$ are Lindelof since by Claim 2.14 each is a countable union of Lindelof subspaces. And by 2.10, $L$ has the covering property. ∎
Now it suffices to show $K$ has the covering property.
Claim 2.16.
$K$ has the covering property
Proof.
Note that $K$ is now a graph of some strictly decreasing function $g$. By the definition of $E$, such $g$ can only be defined on a countable set, hence the graph is countable hence Lindelof. ∎
This shows that each $A_{q}$ has the covering property completing the proof that $A_{1}$ has the covering property. ∎
Now with all the observations and claims, we obtain a countable cover of $A$ witnessing that $A$ has the covering property and so by the shoelace lemma, $A$ and $B$ can be separated by open sets, hence $E^{2}$ is normal. ∎
Now one may try to categorize all sets $E\subseteq\mathbb{R}$ such that $(E[\leq])^{2}$ into either $Q$-set or a $2$-entangled set. However, the following example shows that is not the case.
@@
Let $f:\mathbb{R}\to\mathbb{R}$ be any one-to-one increasing function such that $dom(f)$ is of cardinality $\aleph_{1}$ and $\{x\in dom(f):f(x)\neq x\}$ is of cardinality $\aleph_{1}$. Wlog, we can assume $f(x)\neq x$ for all $x\in dom(f)$. We will preserve this function $f$ while making it concentrated on rationals and killing all the strictly decreasing functions on uncountable sets. Let $G_{f}={(u_{\alpha},v_{\alpha}):\alpha<\omega_{1}}$. As in 2.8, let $\mathcal{F}=\{f_{\alpha}:\alpha<\omega_{1}\}$ be the set of strictly decreasing functions, $\mathcal{A}=\{A_{\alpha}:\alpha<\omega_{1}\}$ be the set of open sets containning $\mathbb{Q}$ and for each $\alpha$, $C_{\alpha}=\cap_{\beta\leq\alpha}A_{\alpha}\setminus\mathbb{Q}$. We now define elements $x_{\alpha}\in\mathbb{R}$ recursively as follows:
Stage 0: Take $x_{0}\in C_{0}\setminus\{u_{0},v_{0}\}$ such that $f_{0}(x_{0})\not=x_{0}$
@@
$X_{\alpha}$ $=\{x_{\beta}:\beta<\alpha\}\cup\{u_{\beta}:\beta<\alpha\}\cup\{v_{\beta}:\beta<\alpha\}$
@@
Note that all of the above sets are countable, and $C_{\alpha}$ is uncountable. Hence, we can choose $x_{\alpha}\in C_{\alpha}\setminus(X_{\alpha}\cup I_{\alpha}\cup P_{\alpha})$, such that either $x_{\alpha}\not\in dom(f_{\alpha})$ or $f_{\alpha}(X_{\alpha})\not\ni x_{\alpha}$.
Finally we define $E=\mathbb{Q}\cup\{x_{\alpha}:\alpha<\omega_{1}\}\cup\{u_{\alpha}:\alpha<\omega_{1}\}\cup\{v_{\alpha}:\alpha<\omega_{1}\}$.
Now $E$ is not entangled since the uncountable one-to-one increasing function $f$ exists. $(2)$ and $(3)$ follows exactly as in 2.8. ∎
@@ 3. $\lambda$-sets and the Sorgenfrey square
If $X$ is a $\lambda$-set, then $(X[\leq])^{2}$ is pseudonormal.
@@
Let $A,B$ be closed sets and assume $B$ is countable. We will again use shoelace lemma, and so suffices to show that $A$ has the covering property with respect to $B$. Since $B$ is countable it clearly has the covering property with respect to $A$. We begin by defining the following sets:
@@
Note that $B_{0},B_{1}$ are countable, so $A_{0}$ and $A_{1}$ are Lindelöf. Hence $A_{0}$ and $A_{1}$ have the covering property. Now it suffices to prove
@@
We know that there are families $\{F_{n}:n\in\omega\}$ and $\{H_{n}:n\in\omega\}$ of relatively closed (in the Eucliden topology) subsets of $X$ such that
@@
We see $W\subseteq\cup_{n,m}W_{n,m}$ and each $W_{n,m}$ is Sorgenfrey open. We know that $W_{n,m}\cap B=\emptyset$, so it suffices to show $(\overline{W_{n,m}}\setminus W_{n,m})\cap B=\emptyset$.
@@
From the previous example, we know that it is possible for $(X[\leq])^{2}$ to be normal, while $X$ is not a $\lambda$-set. However, we do have the following partial converse analogous to Proposition 2.4.
@@
Since $f$ is a homeomorphism, the graph $G_{f}$ is closed and discrete, any subset of $G_{f}$ is closed in $X^{2}$. Let $A$ be any countable subset, we have $G_{f|_{A}}$ is countable as well, hence $G_{f|_{A}}$ and $G_{f|_{X\setminus A}}$ are closed and can be separated by open sets. Then by 2.2, $X\setminus A$ is relatively $F_{\sigma}$ in $X$, hence $X$ is a $\lambda$-set. ∎
@@
Is it relatively consistent with ZFC that for any uncountable set of reals $X$, $(X[\leq])^{2}$ is not normal?

### v2

@@ 2. Q-sets and the Sorgenfrey square
Let $E$ be any subset of $\mathbb{R}$} that admits no uncountable strictly decreasing functions (e.g., the set constructed in Theorem 2.9), then $(E[\leq])^{2}$ is normal.
@@
Following the notation of 2.10 (1): $G_{q}$ is the graph of a strictly decreasing function, but by hypothesis on $E$, each $G_{q}$ is countable, hence Lindelöf. By 2.10 (2), $(E[\leq])^{2}$ is normal. ∎
Now one may ask if every $E\subseteq\mathbb{R}$ such that $(E[\leq])^{2}$ is normal is either a $Q$-set or is $2$-entangled. However, the following example shows that is not the case.
@@
We will construct our set $E$ by preserving the strictly increasing function $x\mapsto x+1$ while at the same time making it concentrated on $\mathbb{Q}$} and killing all the strictly decreasing functions on uncountable sets. As in 2.9, let $\mathcal{F}=\{f_{\alpha}:\alpha<\omega_{1}\}$ enumerate all strictly decreasing functions defined on Borel subsets of the reals, $\mathcal{A}=\{A_{\alpha}:\alpha<\omega_{1}\}$ an enumeration of all Euclidean open sets containing $\mathbb{Q}$ and, for each $\alpha<\omega_{1}$, let $C_{\alpha}=\cap_{\beta\leq\alpha}A_{\alpha}\setminus\mathbb{Q}$. We now define elements $x_{\alpha}\in\mathbb{R}$ recursively as follows:
Stage 0: Take $x_{0}\in C_{0}$ such that $f_{0}(x_{0})\not=x_{0}$ and $x_{0}\not\in f_{0}(\mathbb{Q})\cup f_{0}^{-1}(\mathbb{Q})$
@@
$X_{\alpha}$ $=\{x_{\beta}:\beta<\alpha\}\cup\{x_{\beta}+1:\beta<\alpha\}\cup\mathbb{Q}$
@@
Note that all of the above sets are countable, and $C_{\alpha}$ is uncountable. Hence, we can choose $x_{\alpha}\in C_{\alpha}\setminus(X_{\alpha}\cup I_{\alpha}\cup P_{\alpha})$.
Finally we define $E=\mathbb{Q}\cup\{x_{\alpha}:\alpha<\omega_{1}\}\cup\{x_{\alpha}+1:\alpha<\omega_{1}\}$.
Verify Property (1): $E$ is not 2-entangled since the uncountable one-to-one increasing function $x\mapsto x+1$ exists.
Verify Property (2): This follows exactly as in 2.9.
Verify Property (3): We will show $(E[\leq])^{2}$ is normal using 2.11. Fix any strictly decreasing function $g$ defined on subsets of $E$. By 2.7, there exists $F$ a strictly decreasing homeomorphism defined from a subset of $E$ into $E$ such that $dom(F)\setminus dom(g)$ is countable. There exists an $\alpha_{0}<\omega_{1}$ such that $F=f_{\alpha_{0}}$ on $dom(F)$. We will show the $f_{\alpha_{0}}$ can only be defined on a countable set, hence $dom(g)$ is countable and $E$ admits no strictly decreasing function on uncountable subsets. Indeed, for all $\beta>\alpha_{0}$, $f_{\alpha_{0}}(x_{\beta})$ can not be mapped to $\{x_{\beta}:\beta\geq\alpha_{0}\}\}$. For all $\gamma\geq\alpha_{0}$,
[…]
$x_{\beta}\not\in P_{\beta}\ni f_{\alpha_{0}}^{-1}(x_{\gamma})\implies f_{\alpha_{0}}(x_{\beta})\neq x_{\gamma}$
[…]
$x_{\gamma}\not\in P_{\gamma}\ni f(x_{\beta})\implies x_{\gamma}\neq f_{\alpha_{0}}(x_{\beta})$
For the same reason, $f_{\alpha_{0}}(x_{\beta})$ cannot be mapped to $\{x_{\beta}+1:\beta\geq a_{0}\}$. Since $f_{\alpha_{0}}$ is one-to-one, $f_{\alpha_{0}}$ has countable domain. ∎
@@ 3. $\lambda$-sets and the Sorgenfrey square
If $X$ is a $\lambda$-set, then $(X[\leq])^{2}$ is pseudo-normal.
@@
Let $A,B$ be closed sets and assume $B$ is countable. Recall the shoelace lemma: to separate $A$ and $B$ it suffices to cover each set by a countable family of open sets whose closures are disjoint from the other closed set.
It suffices to show that $A$ has the covering property with respect to $B$. Since $B$ is countable it clearly has the covering property with respect to $A$. We begin by defining the following sets:
@@
Note that $B_{0}$ is countable, then $p_{0}^{-1}(B_{0})=\bigcup_{b\in B_{0}}p_{0}^{-1}(b)$ is a countable union of Lindelöf spaces which is still Lindelöf. Since $A$ is Sorgenfrey closed, $A_{0}$ is a closed subspace of a Lindelöf space, which is still Lindelöf. By the similar reason, $A_{1}$ is Lindelöf as well. Hence both $A_{0}$ and $A_{1}$ have the covering property. Now it suffices to prove
@@
We know that there are families $\{F_{n}:n\in\omega\}$ and $\{H_{n}:n\in\omega\}$ of relatively closed (in the Euclidean topology) subsets of $X$ such that
@@
We see $W\subseteq\cup_{n,m}W_{n,m}$ and each $W_{n,m}$ is Sorgenfrey open. We know that $W_{n,m}\cap B=\emptyset$, so it suffices to show $(\overline{W_{n,m}}^{\sigma}\setminus W_{n,m})\cap B=\emptyset$.
@@
From the previous example, we know that $(X[\leq])^{2}$ may be normal, while $X$ is not a $\lambda$-set. However, we do have the following partial converse analogous to Proposition 2.4.
@@
Since $f$ is strictly decreasing we have that the graph $G_{f}$ is closed and discrete in $(X[\leq])^{2}$. Therefore any subset of $G_{f}$ is closed. Let $A$ be any countable subset, we have $G_{f|_{A}}$ is countable as well, hence $G_{f|_{A}}$ and $G_{f|_{X\setminus A}}$ are closed and can be separated by Sorgenfrey open sets. Then by 2.2, $X\setminus A$ is relatively $F_{\sigma}$ in $X$, hence $X$ is a $\lambda$-set. ∎
@@
Is it relatively consistent with ZFC that $(X[\leq])^{2}$ is never normal for $X$ an uncountable set of reals?